OpenStudy (anonymous):

sol e the equation below for x 4x-5y=20

OpenStudy (anonymous):

isolate the x add 5y to both sides 4x=20+5y divide by 4 x=5+(5/4)y

OpenStudy (anonymous):

ok thank you so very much i have a question its a different problem